### v1

@@ Abstract.
We study one– and two–dimensional periodic tight-binding models under the presence of a potential that grows to infinity in one direction, hence preventing the particles to escape in this direction (the soft wall). We prove that a spectral flow appears in these corresponding edge models, as the wall is shifted. We identity this flow as a number of Bloch bands, and provide a lower bound for the number of edge states appearing in such models.
© 2024 by the authors. This paper may be reproduced, in its entirety, for non-commercial purposes.
@@ 1. Introduction
The goal of this paper is to study edge states in terminated periodic structures described by tight–binding (TB) Hamiltonians. Such TB operators are extensively studied in condensed matter, as they provide simple models which correctly reproduce the physics of more complex continuous ones (represented e.g. by continuous Schrödinger operators). This research is motivated by models from solid state physics such as the one-dimensional Su-Schrieffer-Heeger (SSH) chain [30] and the two-dimensional Wallace model for graphene [32]. Indeed, it is known for these models that edge states may appear when these periodic systems are restricted to a halfspace. For the Wallace model for instance, the presence of these edge states depends strongly on the direction of the cut. Notable examples are the zigzag cut, where a flat band of edge modes appear, and the armchair cut, where no edge states appear, see [1, 25, 7, 11, 9, 10]. Even in the simple one–dimensional SSH model, an edge stat appears or disappears depending on which of the nonequivalent bonds is cut. In these one-dimensional systems, the topological versus trivial nature of experimentally observed edge states is still a topic of hot debate (see e.g. [26] for theoretical considerations and [5] for recent experimental results) ever since the first reports of Majorana fermions in superconducting wires [6, 24].
In the present article, we give a framework to study the appearance of edge modes in general TB models arising in a wide variety of physical situations, such as graphene sheets, superconducting wires or optical lattices. However, instead of working with hard truncations (Dirichlet boundary conditions), as it is usually done in numerical simulations, we are interested in edge modes that arise from the lattice termination by a soft wall, that is, a continuous potential barrier that impedes the propagation in one halfspace. The main reason is that one cannot translate properly a hard cut in TB models. To illustrate the problem, imagine a TB problem on the line $\mathbb{Z}$, with the presence of a hard wall whose position is parametrized by $t\in\mathbb{R}$, with the convention that one ¡¡deletes¿¿ all sites on the left of $t$. Then, when $t$ varies in an interval of the form $(n-1,n)$, nothing changes, but when $t$ crosses the site $n$, the corresponding Hilbert space suddenly switches from $\ell^{2}(\{n,n+1,\cdots\})$ to $\ell^{2}(\{n+1,n+2,\cdots\})$. The corresponding operators are not continuous in $t$, which makes this case difficult to study. Note that, for some experimental realizations of periodic systems, confinement by a smooth potential instead of a sharp lattice termination is easily achieved. Numerical results on the effect of soft confinement in different configurations can be found for instance in [4, 14, 15].
@@
When studying the analogue of this problem for continuous models, that is for Schrödinger operators with periodic potentials in $\mathbb{R}$ or $\mathbb{R}^{2}$, one can use the spectral flow to study the effect of hard truncations, modelled by a Dirichlet boundary condition along a line parametrized by a translation parameter $t$ (the cut). These models have been studied in [16, 17, 18], where it is proved that a spectral flow appears as the cut is translated. In addition, this spectral flow can be computed explicitly, and is related to a number of Bloch bands of the initial bulk operator. The results of the present paper are the discrete analogue of these, although the techniques of proof are rather different.
@@
Let us give a short description of our main result in the one–dimensional setting. We consider a general TB bulk operator $H$ acting on $L^{2}(\mathbb{Z},\mathbb{C}^{N})$, of the convolution form
@@
where for each $n\in\mathbb{Z}$, $h(n)$ is an $N\times N$ matrix such that $h^{*}(n)=h(-n)$. We assume throughout that $\sum_{n}\left\|h_{n}\right\|_{\rm op}<\infty$, which guarantees the boundedness of the operator by Young’s inequality (see Lemma 2.1 below). By taking a Fourier transform, the spectrum of $H$ is purely essential, of the form
@@
$\forall\Psi\in\ell^{2}(\mathbb{Z},\mathbb{C}^{N}),\qquad(W(t)\Psi)_{n}:=w(n-t)\Psi_{n}.$
@@
It is not difficult to see that the family $t\mapsto H^{\sharp}(t)$ is $1$–quasi–periodic, in the sense that $H^{\sharp}(t+1)=\tau_{1}^{*}H^{\sharp}(t)\tau_{1}$, where $\tau_{1}$ is the usual translation operator $(\tau_{1}\Psi)_{n}:=\Psi_{n-1}$. We prove in Proposition 2.4 below that for all $t\in\mathbb{R}$, $H^{\sharp}(t)$ is self-adjoint with a domain independent of $t$, and that the essential spectrum of $H^{\sharp}(t)$ is also independent of $t$. However, some extra eigenvalues may appear in the essential gaps as $t$ moves. For $E\in\mathbb{R}\setminus\sigma_{\rm{bulk}}$, we denote by
@@
In particular, the spectral flow is independent of the precise expression for the soft-wall. A nonzero spectral flow implies that, at least for some values of $t\in\mathbb{R}$, the set $\sigma_{\rm{edge}}(t):=\sigma(H^{\sharp}(t))\setminus\sigma_{\rm{bulk}}$, sometimes called the edge spectrum, is non empty. It only consists of eigenvalues. The corresponding eigenvectors are called edge states, and describe localized modes which are blocked by the wall on the left, and which cannot propagate in the medium on the right.
@@
Let $w:\mathbb{R}\mapsto{\mathcal{S}}_{N}$ be a soft–wall which is $\nu$–Lipschitz, for some $\nu>0$. If $E$ belongs to an essential gap of $H$, then for all $t_{0}\in\mathbb{R}$ the operator $H^{\sharp}(t_{0})$ has at least ${\mathcal{N}}(E)$ eigenvalues in each interval of the form $(\lambda,\lambda+\nu]$ in this gap.
In other words, the density of eigenvalues in the gap containing $E$ is ${\mathcal{N}}(E)/\nu$ If $\nu$ is small, the potential $W$ increases slowly to +$\infty$+ ∞ on the left. It somehow acts as a local energy shift. It is therefore not so surprising that many eigenvalues appear in all gaps in this situation. On the other hand, if $\nu$ is very large, (imagine a situation where one approximates a hard cut by a sequence of soft wall with $\nu\to\infty$) then the spectral flow does not give information on the spectrum of $H^{\sharp}(t_{0})$: it may or may not contain extra eigenvalues. Figure 3 below illustrates this situation for the SSH model.
Before we go on, let us recall that there are several ways to derive TB models from continuous ones. One approach, called the strong binding limit, is given in [12, 13, 28], and consists in studying the limit of continuous Schrödinger operators with a periodic potential, when the wells of the potential approach a delta-function centered on each atom. In this limit (similar to the semi-classical one), the wave-function $\Psi$ is replaced by a discrete function giving the amplitude of $\Psi$ associated to each atom. Another way to derive TB models is through the use of Wannier functions. In this case, one finds a countable set of functions $(w_{i})$ which are well-localized around the atoms $({\mathbf{x}}_{i})$, and which span the lowest part of the spectrum of the continuous Hamiltonian. A continuous wave–function $\Psi(x)$ is then approximated by the sum $\sum c_{i}w_{i}(x)$. Thanks to the localization of $w_{i}$, the value $c_{i}w_{i}({\mathbf{x}}_{i})$ can be interpreted as the value $\Psi({\mathbf{x}}_{i})$. The hopping transition between atoms $i$ and $j$ are then $t_{ij}=\langle w_{i},Hw_{j}\rangle$. We refer to [29, 20] for more details.
@@
In Section 4, we explain how to extend our result in the two–dimensional setting. The main observation is that a 2d model, even cut by a (commensurate) soft-wall, is still periodic in the direction along the wall. After a Fourier transform in this direction, we are left with a family of one-dimensional models, indexed by the momentum in the direction along the wall. We prove in Theorem 4.5 that the greater the incommensurability of the cut, the more abundant the edge states. In Section 4.3, we specify to the Wallace model for graphene to illustrate the concepts and give numerical results on the edge states.
@@ 2.1.1. Bulk periodic Hamiltonians
$\forall\Psi\in\ell^{2}(\mathbb{Z},\mathbb{C}^{N}),\quad\|H\Psi\|_{\ell^{2}}=\|h*\psi\|_{\ell^{2}}\leq\|h\|_{\ell^{1}}\|\psi\|_{\ell^{2}},\quad\text{with}\quad\|h\|_{\ell^{1}}:=\sum_{n=-\infty}^{\infty}\|h(n))\|_{{\rm{op}}}.$
@@
Throughout this section, we will use the term periodic to refer to 1-periodic models. If a Hamiltonian is periodic with period $K$ for some integer $K>1$, we can identify $\ell^{2}(\mathbb{Z},\mathbb{C}^{N})$ with $\ell^{2}(K\mathbb{Z},\mathbb{C}^{K\,N})$ hence with $\ell^{2}(\mathbb{Z},\mathbb{C}^{K\,N})$, and obtain a 1-periodic model with larger blocks. See Subsection 2.4 for an example of this.
@@ 2.1.2. Soft wall operators
Our goal is to study such periodic Hamiltonians, in the presence of a soft wall. Let us precise the class of walls that we use in this article.
@@
[…]
@@
Note that $\left\|W(t)-W(0)\right\|_{{\rm{op}}}=\sup_{n\in\mathbb{Z}}\left\|w(n-t)-w(t)\right\|_{{\rm{op}}}\leq\nu|t|$, so this domain is independent of $t$ and we will denote it simply by ${\mathcal{D}}(W)$. Let $\mu_{1,x}\leq\cdots\leq\mu_{N,x}$ be the eigenvalues of $w(x)$ for all $x\in\mathbb{R}$. Since $w$ is $\nu$–Lipschitz, all the curves $x\mapsto\mu_{1,x}$ are $\nu$–Lipschitz as well. In addition, all these curves go from +$\infty$+ ∞ to 00 on $\mathbb{R}$. As $W(t)$ is block diagonal with block elements $w(n-t)$, we directly deduce that
@@
The spectrum is purely discrete, composed of eigenvalues. They are all of finite multiplicities (except maybe 00), and the only accumulation point is 00. In addition, the map $t\mapsto W(t)$ is $1$–quasi–periodic, in the sense that
@@
In particular, the spectrum is $1$–periodic (which can be read directly from (6)). See also Fig. 1 for an illustration.
@@
We define the edge operator $H^{\sharp}(t):=H+W(t)$. Let us record some properties for this operator.
@@
    The map $t\mapsto H^{\sharp}(t)$ is norm–resolvent continuous and $1$–quasi–periodic, in the sense that $H^{\sharp}(t+1)=\tau_{1}^{*}H^{\sharp}(t)\tau_{1}$, where $\tau_{1}$ is the usual translation operator $(\tau_{1}\Psi)_{n}:=\Psi_{n-1}$;
@@
The quasi-periodicity follows from the quasi-periodicity of $W(t)$ and the periodicity of $H$. Since $\tau_{1}$ is unitary, it implies that the spectrum of $H^{\sharp}(t)$ is 1-periodic in $t$.
@@
$H^{\sharp}=H^{\sharp}_{L}\oplus H^{\sharp}_{R}+K,\quad\text{with}\quad H^{\sharp}_{L}=\Pi_{L}H^{\sharp}\Pi_{L}^{*},\quad H^{\sharp}_{R}(t)=\Pi_{R}H^{\sharp}\Pi_{R}^{*}$
@@
Since $h_{\ell}$ converges to $h$ is $\ell^{1}(\mathbb{Z},\mathbb{C}^{N\times N})$, we have $\|H-H_{\ell}\|_{\rm op}\to 0$ as $\ell\to\infty$ by Young’s inequality (see the proof of Lemma 2.1). So we also have $\|K-K_{\ell}\|_{\rm op}\to 0$, with $K_{\ell}:=\Pi_{R}H_{\ell}\Pi_{L}^{*}+\Pi_{L}H_{\ell}\Pi_{R}^{*}$. In addition, $K_{\ell}$ is finite rank for all $\ell>0$, so $K$ is compact, as stated.
@@
A convenient way to identify $\sigma_{{\rm{ess}}}(H_{R})$ is to note that, again by compactness of $K$,
$\sigma_{\rm{bulk}}=\sigma_{\rm{ess}}(H_{L})\cup\sigma_{\rm{ess}}(H_{R}),$
but $\sigma_{\rm{ess}}(H_{L})=\sigma_{\rm{ess}}(H_{R})$, since a reflection around $-1/2$ maps $H_{L}$ to $H_{R}$. Hence, $\sigma_{{\rm{ess}}}(H^{\sharp}_{R})=\sigma_{\rm{bulk}}$.

### v3

@@ Abstract.
We study one– and two–dimensional periodic tight-binding models in the presence of a potential that grows to infinity in one direction, hence preventing the particles to escape in this direction (the soft wall). We prove that a spectral flow appears in these edge models, as the wall is shifted with respect to the lattice. We identity this flow with the number of Bloch bands. This provides a lower bound for the number of edge states appearing in such models. For the two-dimensional case, we compute the spectral flow for edges that have any rational orientation with respect to the lattice. The results are illustrated by applying them to the one-dimensional SSH chain and the Wallace model for graphene.
© 2025 by the authors. This paper may be reproduced, in its entirety, for non-commercial purposes.
@@ 1. Introduction
The goal of this paper is to study edge states in terminated periodic structures described by tight–binding (TB) Hamiltonians. Such TB operators are extensively studied in condensed matter, as they provide simple models which correctly reproduce the physics of more complex ones (represented e.g. by Schrödinger operators acting on the continuum). This research is motivated by models from solid state physics such as the one-dimensional Su-Schrieffer-Heeger (SSH) chain [34] and the two-dimensional Wallace model for graphene [36]. Indeed, it is known for these models that edge states may appear when these periodic systems are restricted to a halfspace. For the Wallace model for instance, the presence of these edge states depends strongly on the direction of the cut. Notable examples are the zigzag cut, where a flat band of edge modes appear, and the armchair cut, where no edge states appear, see [1, 29, 8, 13, 11, 12]. Even in the simple one–dimensional SSH model, an edge state appears or disappears depending on which of the nonequivalent bonds is cut. In these one-dimensional systems, the topological versus trivial nature of experimentally observed edge states is still a topic of hot debate (see e.g. [30] for theoretical considerations and [5] for recent experimental results) ever since the first reports of Majorana fermions in superconducting wires [6, 28].
In the present article, we give a framework to study the appearance of edge modes in general TB models arising in a wide variety of physical situations, such as graphene sheets, superconducting wires or optical lattices. However, instead of working with hard truncations (Dirichlet boundary conditions), as it is usually done in numerical simulations, we are interested in edge modes that arise from the lattice termination by a soft wall, that is, a continuous potential barrier that impedes the propagation in one halfspace. The main reason is that one cannot translate properly a hard cut in TB models. To illustrate the problem, imagine a TB problem on the line $\mathbb{Z}$, with the presence of a hard wall whose position is parametrized by $t\in\mathbb{R}$, with the convention that one deletes all sites on the left of $t$. Then, when $t$ varies in an interval of the form $(n-1,n)$, nothing changes, but when $t$ crosses the site $n$, the corresponding Hilbert space suddenly switches from $\ell^{2}(\{n,n+1,\cdots\})$ to $\ell^{2}(\{n+1,n+2,\cdots\})$. The corresponding operators are not continuous in $t$, which makes this case difficult to study. Note that, for some experimental realizations of periodic systems, confinement by a smooth potential instead of a sharp lattice termination is easily achieved. Numerical results on the effect of soft confinement in different configurations can be found for instance in [4, 16, 17].
@@
When studying the analogue of this problem for models set on the continuum, that is for Schrödinger operators with periodic potentials in $\mathbb{R}$ or $\mathbb{R}^{2}$, one can use the spectral flow to study the effect of hard truncations, modelled by a Dirichlet boundary condition along a line parametrized by a translation parameter $t$ (the cut). These models have been studied in [24, 23, 25, 26, 10, 18, 19, 20], where it is proved that a spectral flow appears as the cut is translated. In addition, this spectral flow can be computed explicitly, and is related to a number of Bloch bands of the initial bulk operator. The results of the present paper are the discrete analogue of these, although the techniques of proof are rather different.
@@
Let us give a short description of our main result in the one–dimensional setting. We consider a general TB bulk operator $H$ acting on $\ell^{2}(\mathbb{Z},\mathbb{C}^{N})$, of the convolution form
@@
where for each $n\in\mathbb{Z}$, $h(n)$ is an $N\times N$ matrix such that $h^{*}(n)=h(-n)$. We assume throughout that $\sum_{n}\left\|h_{n}\right\|_{\rm op}<\infty$, which guarantees the boundedness of the operator by Young’s inequality (see Lemma 2.1 below). Strictly speaking, the model is set on $a_{0}\mathbb{Z}$ with $a_{0}>0$ the length of the unit cell. In what follows, we take $a_{0}=1$. By taking a Fourier transform, the spectrum of $H$ is purely essential, of the form
@@
$\forall\Psi\in\ell^{2}(\mathbb{Z},\mathbb{C}^{N}),\qquad(W(t)\Psi)_{n}:=w(n-t)\Psi_{n}.$ (2)
@@
It is not difficult to see that the family $t\mapsto H^{\sharp}(t)$ is translation equivariant, in the sense that $H^{\sharp}(t+1)=\tau_{1}^{*}H^{\sharp}(t)\tau_{1}$, where $\tau_{1}$ is the usual translation operator $(\tau_{1}\Psi)_{n}:=\Psi_{n-1}$. We prove in Proposition 2.4 below that for all $t\in\mathbb{R}$, $H^{\sharp}(t)$ is self-adjoint with a domain independent of $t$, and that the essential spectrum of $H^{\sharp}(t)$ is also independent of $t$. However, some extra eigenvalues may appear in the essential gaps as $t$ moves. For $E\in\mathbb{R}\setminus\sigma_{\rm{bulk}}$, we denote by
@@
In particular, the spectral flow is independent of the precise expression for the soft wall. A nonzero spectral flow implies that, at least for some values of $t\in\mathbb{R}$, the set $\sigma_{\rm{edge}}(t):=\sigma(H^{\sharp}(t))\setminus\sigma_{\rm{bulk}}$, sometimes called the edge spectrum, is non empty. It only consists of eigenvalues. The corresponding eigenvectors are called edge states, and describe localized modes which are blocked by the wall on the left, and which cannot propagate in the medium on the right.
Let us briefly explain the idea of the proof with the help of the graphical representation in Figure 1. First, we show that the spectral flow for any soft wall is equal to the spectral flow for a simplified steep wall model that essentially behaves as a step potential with value $\Sigma\gg E$ (Figure 1.1). Next, we show that the spectral flow decouples in a left and right contribution by summing a compact perturbation that erases some bonds (Figure 1.2). Here, the right-hand side of the picture does not contribute to the spectral flow at energy $E$. A similar procedure can be performed on the dislocated model shown in Figure 1.3. In this model, as $t$ increases by one, one site decouples from the otherwise periodic chain. A suitable compact perturbation yields a cut dislocated model shown in Figure 1.4. As before, the left-hand-side does not contribute to the spectral flow while the right-hand-sides coincides with that of the steep wall model. This allows us to show that the spectral flow for any soft wall model equals the spectral flow of the related dislocated model. Then, the spectral flow for the dislocated model can be computed exactly, by interpreting it as the limit of finite chains with periodic boundary conditions.
Figure 1. Schematic illustration of the models in the proof. In the first figure, as $t$ varies from 00 to $1$, the wall is pushed one cell to the right. We prove that all spectral flows coincide for these four scenarios, and explicitly compute the third one.
@@
Let $w:\mathbb{R}\mapsto{\mathcal{S}}_{N}$ be a soft wall which is $\nu$–Lipschitz, for some $\nu>0$. If $E$ belongs to an essential gap of $H$, then for all $t_{0}\in\mathbb{R}$ the operator $H^{\sharp}(t_{0})$ has at least ${\mathcal{N}}(E)$ eigenvalues in each interval of the form $(\lambda,\lambda+\nu]$ in this gap.
Remark 1.4.
Recall that we took the length scale of the chain to be $a_{0}=1$, so $\lambda$ and $\nu$ have the same units. To restore a length scale and work instead on $a_{0}\mathbb{Z}$, it is convenient to define the operator $W$ in (2) by $(W(t)\Psi)_{R}:=w(R-ta_{0})\Psi_{R}$ for $R\in a_{0}\mathbb{Z}$, so that $t\mapsto H(t)$ is translation equivariant with period $1$. If $w(\cdot)$ is $\nu$-Lipschitz, then $W(\cdot)$ and $H(\cdot)$ are $\nu a_{0}$–Lipschitz, and there are ${\mathcal{N}}(E)$ eigenvalues in each interval of the form $(\lambda,\lambda+\nu a_{0}]$.
This theorem states that the density of eigenvalues in the gap containing $E$ is at least ${\mathcal{N}}(E)/\nu$. If $\nu$ is small, the potential $W$ increases slowly to +$\infty$+ ∞ on the left. It somehow acts as a local energy shift. It is therefore not so surprising that many eigenvalues appear in all gaps in this situation. On the other hand, if $\nu$ is very large, (imagine a situation where one approximates a hard cut by a sequence of soft wall models with $\nu\to\infty$) then the spectral flow does not give information on the spectrum of $H^{\sharp}(t_{0})$: it may or may not contain extra eigenvalues. Figure 4 below illustrates this situation for the SSH model.
Before we go on, let us recall that there are several ways to derive TB models from the ones set on the continuum. One approach, called the strong binding limit, is given in [14, 15, 32], and consists in studying the limit of Schrödinger operators acting on $L^{2}(\mathbb{R}^{d})$ with a periodic potential, when the wells of the potential approach a delta-function centered on each atom. In this limit (similar to the semi-classical one), the wave-function $\Psi$ is replaced by a discrete function giving the amplitude of $\Psi$ associated to each atom. Another way to derive TB models is through the use of Wannier functions. In this case, one finds a countable set of functions $(w_{i})$ which are well-localized around the atoms $({\mathbf{x}}_{i})$, and which span the lowest part of the spectrum of the Hamiltonian. A wave–function $\Psi(x)$ is then approximated by the sum $\sum c_{i}w_{i}(x)$. Thanks to the localization of $w_{i}$, the value $c_{i}w_{i}({\mathbf{x}}_{i})$ can be interpreted as the value $\Psi({\mathbf{x}}_{i})$. The hopping transition between atoms $i$ and $j$ are then $t_{ij}=\langle w_{i},Hw_{j}\rangle$. We refer to [33, 22] for more details.
@@
In Section 4, we explain how to extend our result in the two–dimensional setting. The main observation is that a 2d model, even cut by a (commensurate) soft wall, is still periodic in the direction along the wall. After a Fourier transform in this direction, we are left with a family of one-dimensional models, indexed by the momentum in the direction along the wall. We prove in Theorem 4.5 that the greater the incommensurability of the cut, the more abundant the edge states. In Section 4.3, we specify to the Wallace model for graphene to illustrate the concepts and give numerical results on the edge states.
@@ 2.1.1. Bulk periodic Hamiltonians
$\forall\Psi\in\ell^{2}(\mathbb{Z},\mathbb{C}^{N}),\quad\|H\Psi\|_{\ell^{2}}=\|h*\psi\|_{\ell^{2}}\leq\|h\|_{\ell^{1}}\|\psi\|_{\ell^{2}},\quad\text{with}\quad\|h\|_{\ell^{1}}:=\sum_{n=-\infty}^{\infty}\|h(n)\|_{{\rm{op}}}.$
@@
Throughout this section, we will use the term periodic to refer to 1-periodic models. If a Hamiltonian is periodic with period $K$ for some integer $K>1$, we can identify $\ell^{2}(\mathbb{Z},\mathbb{C}^{N})$ with $\ell^{2}(K\mathbb{Z},\mathbb{C}^{K\,N})$ hence with $\ell^{2}(\mathbb{Z},\mathbb{C}^{K\,N})$, and obtain a 1-periodic model with larger blocks. This effectively changes the length–scale $a_{0}$ into $\widetilde{a}_{0}:=Ka_{0}$. See Subsection 2.4 for an example of this.
@@ 2.1.2. Soft wall operators
Our goal is to study such periodic Hamiltonians, in the presence of a soft wall. Let us specify the class of walls that we use in this article.
@@
$\forall v\in\mathbb{C}^{N}\setminus\{0\},\qquad\lim_{x\to-\infty}\langle v,w(x)v\rangle=+\infty\quad\text{and}\quad\lim_{x\to+\infty}\langle v,w(x)v\rangle=0.$
@@
Note that $\left\|W(t)-W(0)\right\|_{{\rm{op}}}=\sup_{n\in\mathbb{Z}}\left\|w(n-t)-w(n)\right\|_{{\rm{op}}}\leq\nu|t|$, so this domain is independent of $t$ and we will denote it simply by ${\mathcal{D}}(W)$. Let $\mu_{1,x}\leq\cdots\leq\mu_{N,x}$ be the eigenvalues of $w(x)$ for all $x\in\mathbb{R}$. Since $w$ is $\nu$–Lipschitz, all the curves $x\mapsto\mu_{1,x}$ are $\nu$–Lipschitz as well. In addition, all these curves go from +$\infty$+ ∞ to 00 on $\mathbb{R}$. As $W(t)$ is block diagonal with block elements $w(n-t)$, we directly deduce that
@@
The spectrum is purely discrete, composed of eigenvalues. They are all of finite multiplicities (except maybe 00), and the only accumulation point is 00. In addition, the map $t\mapsto W(t)$ is translation equivariant, in the sense that
@@
In particular, the spectrum is $1$–periodic (which can be read directly from (7)). Figure 2 illustrates the situation.
@@
We define the edge operator $H^{\sharp}(t):=H+W(t)$. Let us record some properties of this operator.
@@
    The map $t\mapsto H^{\sharp}(t)$ is norm–resolvent continuous and translation equivariant, in the sense that $H^{\sharp}(t+1)=\tau_{1}^{*}H^{\sharp}(t)\tau_{1}$, where $\tau_{1}$ is the usual translation operator $(\tau_{1}\Psi)_{n}:=\Psi_{n-1}$;
@@
The translation equivariance follows from the translation equivariance of $W(t)$ and the periodicity of $H$. Since $\tau_{1}$ is unitary, it implies that the spectrum of $H^{\sharp}(t)$ is 1-periodic in $t$.
@@
$H^{\sharp}=H^{\sharp}_{L}\oplus H^{\sharp}_{R}+K,\quad\text{with}\quad H^{\sharp}_{L}=\Pi_{L}H^{\sharp}\Pi_{L}^{*},\quad H^{\sharp}_{R}=\Pi_{R}H^{\sharp}\Pi_{R}^{*}$
@@
Since $h_{\ell}$ converges to $h$ in $\ell^{1}(\mathbb{Z},\mathbb{C}^{N\times N})$, we have $\|H-H_{\ell}\|_{\rm op}\to 0$ as $\ell\to\infty$ by Young’s inequality (see the proof of Lemma 2.1). So we also have $\|K-K_{\ell}\|_{\rm op}\to 0$, with $K_{\ell}:=\Pi_{R}H_{\ell}\Pi_{L}^{*}+\Pi_{L}H_{\ell}\Pi_{R}^{*}$. In addition, $K_{\ell}$ is finite rank for all $\ell>0$, so $K$ is compact, as stated.
@@
We have $\sigma_{{\rm{ess}}}(H_{R})=\sigma_{\rm{bulk}}$: the inclusion $\sigma_{{\rm{ess}}}(H_{R})\subset\sigma_{\rm{bulk}}$ can be shown directly by extending a singular Weyl sequence for $H_{R}$ by zero to the left. For the opposite inclusion, consider a singular Weyl sequence $(\Psi^{k})_{k\in\mathbb{N}}$ for $H$ and fix a translation $R_{k}$ such that $\left\|\Pi_{L}\Psi^{k}_{R_{k}}\right\|\leq 1/k$, where $\Psi_{R}(n):=\Psi(n-R)$ is the translation of $\Psi$ by $R$. Then $\Pi_{R}\Psi^{k}_{R_{k}}$ is a singular Weyl sequence for $H_{R}$.